### v1

# LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation

Abstract
Conditional layout generation aims to automatically generate visually appealing and semantically coherent layouts from user-defined constraints. While recent methods based on generative models have shown promising results, they typically require substantial amounts of training data or extensive fine-tuning, limiting their versatility and practical applicability. Alternatively, some training-free approaches leveraging in-context learning with Large Language Models (LLMs) have emerged, but they often suffer from limited reasoning capabilities and overly simplistic ranking mechanisms, which restrict their ability to generate consistently high-quality layouts. To this end, we propose LayoutCoT, a novel approach that leverages the reasoning capabilities of LLMs through a combination of Retrieval-Augmented Generation (RAG) and Chain-of-Thought (CoT) techniques. Specifically, LayoutCoT transforms layout representations into a standardized serialized format suitable for processing by LLMs. A Layout-aware RAG is used to facilitate effective retrieval and generate a coarse layout by LLMs. This preliminary layout, together with the selected exemplars, is then fed into a specially designed CoT reasoning module for iterative refinement, significantly enhancing both semantic coherence and visual quality. We conduct extensive experiments on five public datasets spanning three conditional layout generation tasks. Experimental results demonstrate that LayoutCoT achieves state-of-the-art performance without requiring training or fine-tuning. Notably, our CoT reasoning module enables standard LLMs, even those without explicit deep reasoning abilities, to outperform specialized deep-reasoning models such as deepseek-R1, highlighting the potential of our approach in unleashing the deep reasoning capabilities of LLMs for layout generation tasks.
footnotetext: Equal Contribution. Name is ordered by alphabet.
🖂Corresponding Author.
Figure 1: Qualitative evaluation of the multi-stage LayoutCoT framework. Each successive stage systematically refines the spatial arrangement of layout elements, resulting in designs with enhanced coherence and rationality.
1 Introduction
Layout is a fundamental element in graphic design, determined by the deliberate and structured arrangement of visual components. Automated layout generation has recently emerged as an important research topic, aiming to significantly alleviate the workload of designers while addressing diverse user requirements [11, 19, 21, 22, 23, 26]. Current automated layout generation tasks broadly fall into three categories: i) content-aware layouts, which integrate visual or textual content for contextually appropriate designs (e.g., aligning text boxes with salient image regions in posters); ii) constraint-explicit layouts, which strictly adhere to predefined constraints, commonly found in user interface (UI) design and document typesetting; iii) and text-to-layout tasks, which translate abstract textual instructions into spatially coherent layouts, crucial for cross-modal creative tools.
Early approaches [9, 12] mainly employed generative models, learning implicit design rules from extensive labeled datasets. Methods such as LayoutGAN [25] and LayoutDiffusion [43] have shown promising capabilities in generating plausible layouts, yet their reliance on substantial training data and domain-specific fine-tuning limits scalability and versatility. When faced with limited training data, purely generative methods [19, 16, 18, 14, 28] often fail to effectively capture sparse data distributions, restricting generalization and necessitating separate models for distinct tasks. Furthermore, the demand for extensive labeled datasets makes training generative layout models costly and inefficient, highlighting the necessity for a more flexible and versatile layout generation paradigm.
With the advancement of large language models (LLMs) [2, 17, 1, 41, 42], layout generation methods are shifting towards a training-free paradigm. Recent research has shown that LLMs inherently possess a certain understanding of layout principles, such as alignment and spatial coherence [17, 3]. While leveraging this capability, approaches such as LayoutPrompter [28] utilize in-context learning combined with post-processing ranking mechanisms to guide LLMs in layout generation. However, this methodology demonstrates excessive focus on quantitative evaluation metrics, frequently resulting in structurally impractical bounding box configurations containing either disproportionately small elements or physically unfeasible arrangements. Moreover, when confronted with complex layout scenarios, the simplistic reasoning paradigms and elementary ranking systems inherent in such approaches tend to produce suboptimal compositions that violate both human visual perception principles and real-world application constraints.
To unleash the deep reasoning abilities of LLMs in layout generation, we must resolve two key issues: first, how to select suitable contexts as references for layout generation; and second, how to design a chain of thought that enables LLMs to engage in human-like deep reasoning during layout design. Therefore, we propose LayoutCoT, a novel approach that integrates a layout-aware RAG with Chain-of-Thought (CoT) in LLMs to address those above limitations. Specifically, LayoutCoT employs a layout-aware exemplar retrieval mechanism, dynamically selecting suitable reference contexts using layout similarity [32], which better captures structural and semantic similarities compared to previous methods like LayoutPrompter [28]. We further introduce a meticulously designed Chain-of-Thought (CoT) reasoning framework that refines the layout through three sequential stages, each addressing a distinct aspect of the design process. In the first stage, the LLM is prompted to evaluate each element from the specified categories and determine their initial positions, emphasizing logical structure and visual appeal without accounting for interactions among element sizes. In the second stage, the LLM is guided to optimize the sizes and positions of these elements, effectively resolving overlaps or potential crowding issues. Finally, in the third stage, fine-grained mathematical adjustments are applied to specific layout dimensions to avoid unreasonable arrangements—such as the occlusion of salient elements—and to ensure both aesthetic quality and practical viability.
We conducted comprehensive experiments across five public datasets (PKU [14], CGL [46], RICO [29], Publaynet [45], and WebUI [40]), spanning three distinct layout generation tasks. Our results demonstrate that LayoutCoT consistently outperforms existing state-of-the-art methods without requiring any training or fine-tuning. Notably, our CoT module enables standard LLMs, such as GPT-4, to surpass the performance of specialized deep-reasoning models like DeepSeek-R1 [10], underscoring the potential of our approach for practical layout generation tasks. The main contributions of our work are summarized as follows:
  • We introduce LayoutCoT, a method that effectively harnesses the deep reasoning capabilities of general-purpose LLMs through Chain-of-Thought prompting, significantly improving the performance and practicality of training-free layout generation.
  • LayoutCoT requires no task-specific training or fine-tuning, efficiently producing high-quality, visually coherent layouts by decomposing complex generation tasks into clear reasoning steps.
  • Extensive empirical validation demonstrates LayoutCoT’s superior performance and versatility across various tasks and datasets, highlighting its ability to elevate standard LLMs above specialized deep-reasoning counterparts.
2 Related Work
Layout generation has emerged as an active area of research in recent years. It can be broadly divided into three categories: content-aware layout generation [14, 44], constraint-explicit layout generation [21, 22, 23], and text-to-layout generation [15, 27]. Content-aware layout generation aims to place layout elements on a given canvas in a visually pleasing manner while avoiding overlap with any background images on the canvas. Constraint-Explicit layout generation typically involves arranging predefined elements, or elements with specified dimensions, or even specified relationships among elements, on a blank interface in a reasonable configuration. Text-to-layout generation, meanwhile, focuses on creating layouts based on natural language descriptions, which is generally more challenging.
2.1 Generative layout generation
Early work in this field commonly utilized generative models for layout generation. Research based on Generative Adversarial Networks (GANs) [26, 14] and Variational Autoencoders (VAEs) [23, 20] pioneered the application of generative models to layout generation, achieving noticeable success in improving layout aesthetics through latent variable optimization. With the advent of the Transformer [38] architecture, numerous studies [11, 22, 15, 27, 33] emerged that adopted a sequence-to-sequence approach, further enhancing the quality of generated layouts. More recently, the proposal of diffusion [12] models has led to additional methods [16, 18] that bolster generative capabilities in scenarios with explicit layout constraints, offering further advancements in both quality and overall visual appeal. However, the existing generative layout methods are not capable of handling all types of layout generation tasks using a single model. Moreover, these approaches typically rely heavily on extensive datasets and require substantial training or fine-tuning, making practical applications challenging.
2.2 LLM-based Layout Generation
Large language models have demonstrated exceptional few-shot performance [2, 1, 6, 37] across various natural language processing (NLP) tasks. Recent studies have shown that LLMs themselves can generate layouts in structured descriptive languages such as HTML [28, 36]. For instance, LayoutNUWA [36] leverages Code Llama’s [34] code completion for layout generation, while PosterLlama [35] fine-tunes models such as MiniGPT-4 [48] to further enhance code completion capabilities for layout generation tasks. However, these methods rely heavily on sophisticated fine-tuning strategies and are dependent on large amounts of data. In contrast, LayoutPrompter [28] and LayoutGPT [7] utilize GPT-3 [2] via in-context learning to achieve training-free layout generation, revealing that LLMs inherently possess the capacity for layout generation. Although these methods do not require extensive, painstaking training and can generalize across various tasks, the quality of their generated layouts remains a concern.
2.3 Retrieval-Augmented Generation and Chain of Thought
Retrieval-Augmented Generation (RAG) [24] is a cutting-edge technique that combines retrieval and generation to enhance the model’s generative capabilities. By retrieving relevant examples or information [30, 5], RAG introduces more comprehensive contextual data during the generation process, thereby improving the quality of the outputs. Content-aware graphic layout generation aims to automatically arrange visual elements based on given content, such as e-commerce product images. In this context, the Retrieval-Augmented Layout Transformer (RALF) [13] leverages the retrieval of layout examples most similar to the input image to assist in layout design tasks. However, the RAG in RALF is constrained by its input paradigm, lacking plug-and-play characteristics and performing suboptimally across other tasks. Chain-of-Thought (CoT) [39, 4, 31] is a method that enhances model reasoning capabilities through the simulation of human thought processes. CoT achieves this by progressively solving problems through phased reasoning, offering more interpretable and coherent solutions to complex tasks. Although CoT has demonstrated powerful reasoning advantages across numerous domains, it has not yet been applied to layout design tasks. We attempt to integrate RAG with CoT for layout generation tasks, utilizing RAG to provide rich contextual information, while employing CoT for phased reasoning and optimization. This synergistic approach aims to enhance the quality and coherence of the generated results, combining the strengths of both methodologies effectively.
Figure 2: Overview of the LayoutCoT framework. Our training-free approach initially employs a Layout-aware RAG to prompt the LLM for a coarse layout prediction, establishing a logical and visually coherent arrangement of elements. This is subsequently refined via a multi-stage Chain-of-Thought (CoT) module, which iteratively enhances the layout by resolving spatial conflicts and fine-tuning dimensions. The proposed framework is versatile and applicable to a wide array of layout generation tasks.
3 Methods
Figure. 2 illustrates the overall structure of LayoutCoT, which consists of three main components: the Layout-aware RAG, the layout coarse generator, and the layout CoT module. Specifically, given the labels and bounding boxes (bboxes) of an input example layout, we use the Layout-aware RAG to select the top-k relevant layout examples as prompting examples. Next, the layout coarse generator takes these prompting examples and the given constraints to produce an initial coarse layout. Finally, both the prompting examples and this coarse layout are fed into the layout CoT module for fine-grained layout adjustments. Because the layout coarse generator lacks a detailed reasoning process, relying solely on the prompting examples and constraints may not fully satisfy the design requirements. Therefore, the layout CoT module breaks down the layout task into a multi-stage thought process; each stage refines the layout, ultimately yielding a more precise and aesthetically pleasing result.
3.1 Layout-aware RAG
We employ a Retrieval-Augmented Generation (RAG) mechanism to dynamically select relevant layout examples as prompts, thereby enhancing the flexibility and coherence of generation. Specifically, let =$\mathcal{L}=\bigl{\{}(b_{i},c_{i})\bigr{\}}_{i=1}^{m}$ and ^=$\hat{\mathcal{L}}=\bigl{\{}(\hat{b}_{j},\hat{c}_{j})\bigr{\}}_{j=1}^{n}$ denote two layouts, where each element consists of a bounding box $b_{i}$ (or $\hat{b}_{j}$) and an associated label $c_{i}$ (or $\hat{c}_{j}$). We use a dissimilarity function from LTSim [32] $D(\mathcal{L},\hat{\mathcal{L}})$ based on a soft matching scheme:
$D(\mathcal{L},\hat{\mathcal{L}})\;=\;\min_{\Gamma}\sum_{(i,j)\in\Omega}\Gamma_{i,j}\,\mu\bigl{(}b_{i},\,c_{i},\;\hat{b}_{j},\,\hat{c}_{j}\bigr{)},\quad\text{s.t.}\;\Gamma_{i,j}\geq 0,$ (1)
where $\Gamma_{i,j}$ represents a fractional assignment aligning element $(b_{i},c_{i})$ with $(\hat{b}_{j},\hat{c}_{j})$, and $\mu(\cdot)$ is a cost function that captures positional and label discrepancies. By solving this optimization, we obtain the minimal “transport cost” between the two layouts.
Next, we transform this cost-based measure into a layout-level similarity via an exponential mapping:
$\mathrm{LTSim}(\mathcal{L},\hat{\mathcal{L}})\;=\;\exp\Bigl{(}-\,D(\mathcal{L},\hat{\mathcal{L}})\Bigr{)}.$ (2)
A lower cost $D(\mathcal{L},\hat{\mathcal{L}})$ implies higher layout similarity; we set the scale parameter to $1.0$ for simplicity [32].
To retrieve the top-$K$ examples for a given layout $\mathcal{L}$, we consider a database $\mathbf{DB}=\{\hat{\mathcal{L}}_{1},\hat{\mathcal{L}}_{2},\dots,\hat{\mathcal{L}}_{N}\}$. We compute $\mathrm{LTSim}\bigl{(}\mathcal{L},\hat{\mathcal{L}}_{i}\bigr{)}$ for each candidate ^$\hat{\mathcal{L}}_{i}\in\mathbf{DB}$, and select the top-$K$ layouts that yield the highest scores:
$\mathcal{R}_{K}(\mathcal{L})\;=\;\mathrm{TopK}_{\,\hat{\mathcal{L}}\in\mathbf{DB}}\,\mathrm{LTSim}\bigl{(}\mathcal{L},\,\hat{\mathcal{L}}\bigr{)}.$ (3)
We then collect these ^$\hat{\mathcal{L}}_{k}\in\mathcal{R}_{K}(\mathcal{L})$ as prompting examples in subsequent generation stages. By leveraging the soft matching mechanism in $\mathrm{LTSim}$, the Layout-aware RAG remains robust even when the input and candidate layouts differ in the number or types of elements.
3.2 Coarse Layout Generation
After retrieving the top-$K$ examples via RAG (Section 3.1), we convert each selected layout ^$\hat{\mathcal{L}}_{k}$ into a pre-defined HTML format. Let $\mathcal{R}_{K}(\mathcal{L})=\{\hat{\mathcal{L}}_{1},\dots,\hat{\mathcal{L}}_{K}\}$ be the retrieved set; we encode each ^$\hat{\mathcal{L}}_{k}$ as an HTML snippet and combine all of them with the given constraint $c$. Formally, we provide the large language model (LLM) with:
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\,c\Bigr{\}},$
where $\mathrm{HTML}(\hat{\mathcal{L}}_{k})$ denotes the HTML representation of layout ^$\hat{\mathcal{L}}_{k}$, and $c$ specifies user-defined constraints.
We then prompt the LLM to generate $n$ candidate layouts
$\mathcal{L}^{1}_{HTML}=\text{LLM}\Bigl{(}\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\,c\}\Bigr{)}.$ (4)
By repeating the above generation process $n$ times, we obtain
$\Bigl{\{}\mathcal{L}^{1}_{HTML},\,\dots,\,\mathcal{L}^{n}_{HTML}\Bigr{\}},$
each in HTML format. To select the most promising one from these $n$ candidates, we employ a ranker $\tau(\cdot)$ following LayoutPrompter [28]. Specifically, we compute a layout quality score $\tau\bigl{(}\mathcal{L}^{j}_{HTML}\bigr{)}$ for each generated layout $\mathcal{L}^{j}_{HTML}$ and pick the highest-scoring layout as our coarse result:
[…]
This coarse layout, $\mathcal{L}^{\mathrm{coarse}}_{HTML}$, preserves the key structural constraints specified by the user while serving as the initial configuration for subsequent refinements.
Figure 3: Details of the CoT Module. We illustrate the overall conversational logic of the multi-stage CoT. Depending on the type of task, there are slight variations in the details to adapt to specific requirements.
3.3 Chain of Thought for Layout Refinement
Once we have obtained the coarse layout $\mathcal{L}^{\mathrm{coarse}}_{HTML}$ under constraint $c$, we employ a multi-stage Chain-of-Thought (CoT) strategy to iteratively refine the layout. Concretely, we break the entire layout generation task (still under the user-defined constraint $c$) into $i$ simpler refinement subproblems, denoted by $\{Q_{L}^{1},Q_{L}^{2},\dots,Q_{L}^{i}\}$. Each $Q_{L}^{t}$ ($1\leq t\leq i$) focuses on a specific aspect of improving the layout (e.g., rearranging bounding boxes, reducing overlaps, or adjusting spacing among elements).
Iterative Refinement Procedure. At the first refinement step ($t=1$), we feed the following information to the large language model (LLM):
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;\mathcal{L}^{\mathrm{coarse}}_{HTML},\;c,\;Q_{L}^{1}\Bigr{\}},$
where $\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K})\}$ are the top-$K$ prompting examples (cf. Section 3.1), $\mathcal{L}^{\mathrm{coarse}}_{HTML}$ is the initial coarse layout, $c$ is the user-defined constraint, and $Q_{L}^{1}$ is the first refinement subproblem. The LLM then outputs:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{1}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}^{\mathrm{coarse}}_{HTML},c,\;Q_{L}^{1}\Bigr{)}.$ (6)
At the $t$-th refinement step ($1<t\leq i$), the input is updated with the newly refined layout from the previous step, $\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}}$, together with the same top-$K$ examples, constraint $c$, and the next subproblem $Q_{L}^{t}$:
$\Bigl{\{}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}},\;c,\;Q_{L}^{t}\Bigr{\}},$
yielding:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{t}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}_{HTML}^{\mathrm{refine}_{t-1}},c,\;Q_{L}^{t}\Bigr{)}.$ (7)
After completing all $i$ subproblems, the final refined layout is:
$\mathcal{L}_{HTML}^{\mathrm{refine}_{i}}=\text{LLM}\Bigl{(}\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\dots,\mathcal{L}_{HTML}^{\mathrm{refine}_{i-1}},c,\;Q_{L}^{i}\Bigr{)}.$
This multi-stage, question-and-answer refinement process allows the LLM to tackle each localized challenge in turn, resulting in a more coherent and aesthetically pleasing final layout.
Table 1: Dataset statistics.
Dataset Tasks Training Set Test Set Element Types
PKU content-aware layout generation 9,974 905 3
CGL content-aware layout generation 38,510 1,647 5
RICO constraint-explicit layout generation 31,694 3,729 25
PubLayNet constraint-explicit layout generation 311,397 10,998 5
WebUI text-to-layout 3,835 487 10
Table 2: Content-Aware layout generation task results, all obtained under unconstrained generation conditions. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed, while both LayoutPrompter and LayoutCoT use ’gpt-4’ as the model.
CGL PKU
Method Training-free Content Graphic Content Graphic
[…]
CGL-GAN [47] 0.489 0.268 0.147 0.0420 0.275 0.244 0.269 0.876 - 0.219 0.175 0.226 0.0062 0.575 0.259 0.0605 0.707 -
DS-GAN [14] 0.451 0.224 0.193 0.0485 0.370 0.301 0.075 0.893 - 0.209 0.217 0.229 0.0046 0.585 0.424 0.026 0.903 -
RALF [13] 0.336 0.197 0.247 0.0023 0.943 0.884 0.027 0.995 0.874 0.171 0.150 0.246 0.0007 0.893 0.840 0.018 0.998 0.981
LayoutPrompter [28] 0.415 0.184 0.227 0.0091 0.338 0.317 0.0049 0.996 0.812 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997 0.865
LayoutCoT 0.267 0.182 0.225 0.0018 0.515 0.501 0.0023 0.994 0.803 0.214 0.180 0.264 0.0016 0.925 0.849 0.0004 1.000 0.806
LayoutCoT (ours) 0.170 0.101 0.260 0.0005 0.958 0.958 0.0018 0.998 0.939 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000 1.000
4 Experiments
4.1 Experimental Setup
We conduct detailed experiments on three different types of tasks: content-aware, constraint-explicit, and text-to-layout, using five distinct datasets: PKU [14], CGL [46], RICO [29], Publaynet [45], and WebUI [40]. Since no official test set splits are provided for the PKU and CGL datasets, we used 9,974 labeled images in the PKU dataset as the training set and the database for the Layout-aware RAG module, with 905 unlabeled images serving as the test set. In the CGL dataset, we took 38,510 images as the training set and the database for the Layout-aware RAG module, and selected an additional 1,647 images (with no overlap with the training set) as the test set. For other datasets, we adopt the official data splits, as detailed in Table .1
4.2 Implement Details
In our experiments, we provide details on the choice of the $k$-value for the top-$k$ retrieval in the Layout-aware RAG. When generating the Coarse Layout, we follow the approach of LayoutPrompter [28] and set $k$=10. For the CoT stage, we set $k$=4. We also adopt LayoutPrompter’s setting of $n$=10 for the number of generation times by LLMs when selecting the Coarse Layout. Due to the ’Text-Davinci-003’ model used by LayoutPrompter is no longer available, we replace the LLM with ’gpt-4’ for comparison. Furthermore, when conducting the LayoutCoT experiments, we also use ’gpt-4’ as the LLM. In addition, we will provide the performance results of ’DeepSeek-R1’ after removing the CoT module.
4.3 Evaluation Metrics
For the constraint-explicit layout generation and text-to-layout tasks, we utilize four standard metrics for evaluation: Alignment measures how well the elements within the layout align with each other. Overlap calculates the area of overlap between any two elements in the layout. Maximum IoU, mIoU assesses the maximum intersection over union between the generated layout and the actual layout, providing a measure of layout accuracy. Frechet Inception Distance (FID) quantifies the similarity between the distribution of generated layouts and the distribution of real layouts, using features extracted from a deep neural network trained on images. For the content-aware layout generation task, we reference the eight metrics used in [14] along with a newly designed metric $R_{e}$ that we introduce. The metric $R_{e}$ is used to assess the reasonableness of each element’s size within the layout. Let
$r_{i}\;=\;\frac{S^{i}}{S^{i}_{\text{train}}},\qquad d_{i}\;=\;\bigl{|}\ln r_{i}\bigr{|},\qquad\tau\;=\;\ln(1.1),$
where $S^{i}$ is the predicted average area of $label_{i}$ and $S^{i}_{\text{train}}$ is the corresponding training‑set reference. The constant $\tau$ encodes a ±$\pm 10\%$ tolerance band.
$\operatorname{score}_{i}\;=\;\exp\!\bigl{(}-\max(0,\,d_{i}-\tau)\bigr{)}.$ (8)
Eqa.(8) yields $\operatorname{score}_{i}=1$ whenever $0.9\leq r_{i}\leq 1.1$ and decays exponentially (symmetrically for oversized and undersized elements) once the deviation exceeds 10%.
For $N$ labels, the overall area‑error metric is
$R_{e}^{coarse}\;=\;\frac{1}{N}\sum_{i=1}^{N}\operatorname{score}_{i}.$ (9)
A more extreme‑error‑sensitive alternative is
$R_{e}\;=\;\exp\!\Bigl{(}-\sqrt{\frac{1}{N}\sum_{i=1}^{N}\bigl{(}\max(0,\,d_{i}-\tau)\bigr{)}^{2}}\Bigr{)}.$ (10)
Both $R_{e}^{coarse}$ and $R_{e}$ lie in $(0,1]$.
[…]
4.4 Results on Content-Aware Layout Generation
The quantitative experimental results are shown in Table 2. As seen from the table, our method surpasses all the evaluation metrics used in content-aware layout generation tasks compared to both the training-based baseline methods and another training-free method, LayoutPrompter [28]. Notably, our proposed CoT module enables ”GPT-4” to even outperform the deep reasoning large language model ”DeepSeek-R1” in this layout generation task. Moreover, our performance on the Overlap metric in the PKU dataset was suboptimal, prompting further visual analysis. As illustrated in Figure. 4, the very low scores on the Overlap metric for both LayoutPrompter and LayoutCoT result from numerous bad cases. In these cases, layout bounding boxes are extremely small and densely clustered in the upper-left corner, leading to low Overlap values but highly unaesthetic and impractical layout outcomes. Hence, we introduce the $R_{e}$ metric in Sec. 4.3 to measure the reasonableness of layout sizes, on which we have performed exceptionally well. Qualitative comparative cases are further provided in Figure. 4.
Table 3: Quantitative comparison with other methods on constraint-explicit layout generation tasks. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed.
RICO PubLayNet
Tasks Methods mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$ mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$
Gen-T BLT [22] 0.216 25.633 0.150 0.983 0.140 38.684 0.036 0.196
LayoutFormer++ [19] 0.432 1.096 0.230 0.530 0.348 8.411 0.020 0.008
LayoutPrompter [28] 0.604 2.134 0.0018 0.0285 0.619 1.648 0.0009 0.0037
LayoutCoT 0.718 4.970 0.0015 0.0208 0.795 2.608 0.0001 0.0027
LayoutCoT (Ours) 0.719 2.189 0.0008 0.0220 0.772 1.227 0.0001 0.0024
Gen-TS BLT [22] 0.604 0.951 0.181 0.660 0.428 7.914 0.021 0.419
LayoutFormer++ [19] 0.620 0.757 0.202 0.542 0.471 0.720 0.024 0.037
LayoutPrompter [28] 0.707 3.072 0.0020 0.0505 0.727 1.863 0.0007 0.076
LayoutCoT 0.782 1.252 0.0005 0.0331 0.844 1.701 0.0001 0.034
LayoutCoT (Ours) 0.795 3.506 0.0004 0.0305 0.847 1.430 0.0003 0.004
Gen-R CLG-LO [21] 0.286 8.898 0.311 0.615 0.277 19.738 0.123 0.200
LayoutFormer++ [19] 0.424 5.972 0.332 0.537 0.353 4.954 0.025 0.076
LayoutPrompter [28] 0.583 2.249 0.0007 0.0436 0.627 2.224 0.0008 0.011
LayoutCoT 0.724 2.399 0.0006 0.0301 0.779 1.853 0.0002 0.004
LayoutCoT (Ours) 0.725 3.569 0.0009 0.0145 0.769 1.408 0.0005 0.002
Completion LayoutTransformer [11] 0.363 6.679 0.194 0.478 0.077 14.769 0.019 0.0013
LayoutFormer++ [19] 0.732 4.574 0.077 0.487 0.471 10.251 0.020 0.0022
LayoutPrompter [28] 0.541 2.656 0.0010 0.0405 0.754 1.282 0.0009 0.0008
LayoutCoT 0.682 2.236 0.008 0.0309 0.843 0.840 0.0007 0.0003
LayoutCoT (Ours) 0.716 2.112 0.0008 0.0117 0.770 0.531 4e-5 0.0006
Refinement RUITE [33] 0.811 0.107 0.133 0.483 0.781 0.061 0.029 0.020
LayoutFormer++ [19] 0.816 0.032 0.123 0.489 0.785 0.086 0.024 0.006
LayoutPrompter [28] 0.874 0.225 0.0011 0.1095 0.756 0.143 0.003 0.0006
LayoutCoT 0.869 0.876 0.0006 0.0714 0.801 0.265 0.001 0.0002
LayoutCoT (Ours) 0.890 0.353 0.0003 0.0520 0.839 0.141 0.0004 0.0003
Figure 5: Qualitative Results for the Constraint-Explicit Layout Generation Task. The RICO dataset features a wide variety of label categories, making the completion task on RICO more challenging compared to other task types, we conduct visualizations based on the completion task of the RICO dataset. It is evident that LayoutCoT designs more rational layouts, with better element overlap and alignment compared to other methods.
4.5 Results on Constraint-Explicit Layout Generation
Table 3 presents the results of the quantitative evaluation, where we compare LayoutCoT with several training-based baselines and the training-free method LayoutPrompter. From the metrics, it is evident that, even without any training, LayoutCoT achieves optimal performance on most indicators. The corresponding qualitative results are shown in Figure. 5, which demonstrates that LayoutCoT exhibits outstanding layout control capabilities and generation quality. LayoutCoT not only effectively meets various input constraints—including element type constraints, size constraints, and element relationship constraints—but also produces visually pleasing layouts by avoiding overlap between elements and enhancing overall alignment. These findings collectively validate the effectiveness of LayoutCoT.
4.6 Results on Text-to-Layout
The quantitative metrics are presented in Table 4. Because text-to-layout is one of the most challenging tasks in layout generation, LayoutCoT falls short in the mIoU metric compared to the parse-then-place [27] method. However, it performs exceptionally well on the remaining three metrics. In many scenarios, LayoutCoT more effectively fulfills textual descriptions and excels in aesthetics, avoiding element overlap and maintaining element alignment.
Table 4: Quantitative comparison with baselines on text-to-layout. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed.
Methods mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$
Mockup [15] 0.193 37.012 0.0059 0.4348
parse-then-place [27] 0.684 2.959 0.0008 0.1380
LayoutPrompter 0.174 4.773 0.0009 0.0107
LayoutCoT 0.164 1.018 0.0009 0.0068
LayoutCoT (Ours) 0.199 1.732 0.00006 0.0061
4.7 Ablation Studies
To demonstrate the effectiveness of our approach, we conducted an overall ablation study, as well as ablations focusing on the selection method for prompting examples in the Layout-aware RAG and on the choice of the number of RAG layouts $k$ in the CoT module. All our ablation studies are conducted on the PKU dataset within the representative Content-Aware layout generation task.
Overall ablation study. In Table. 5, we demonstrate the role of each module. It is evident that the Layout-aware RAG and the CoT module complement each other, yielding excellent results across all metrics. Regarding the higher Overlap score when the CoT module is used, we provide an explanation in Sec. 4.4. In reality, although the Overlap metric increases, the CoT module effectively rectifies the layout’s unreasonableness and lack of aesthetic appeal.
Table 5: Abalation study of each method.
[…]
0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997
0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
The Method of the Layout-aware RAG. Table. 6 reports the experimental results obtained with different RAG‑based strategies for selecting prompting examples. Specifically, random denotes randomly selecting $k$ examples; the mIoU strategy chooses examples according to the maximum IoU between saliency maps; DreamSim [8] ranks candidates by their DreamSim scores. The results show that random selection degrades layout‑generation performance—sometimes performing even worse than omitting RAG entirely. By contrast, the mIoU strategy, which assesses relevance between saliency maps, offers a modest improvement, while DreamSim—considering both image and layout coherence—yields further gains over mIoU. Finally, the ltsim method, which measures overall layout similarity, proves the most effective and achieves the best generation results.
Table 6: Abalation study of Layout-aware RAG.
[…]
- 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
random 0.299 0.184 0.241 0.0009 0.808 0.783 0.0023 0.997
mIoU 0.230 0.163 0.265 0.0003 0.951 0.894 0.0010 1.000
DreamSim [8] 0.252 0.086 0.273 0.0004 0.972 0.932 0.0008 0.999
ltsim 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
Selection of $k$. In Table. 7, we present the impact of different numbers of prompting examples on the CoT module’s performance. It can be observed that when $k$=2, the results are relatively poor. This is because having fewer prompting examples means the LLM lacks sufficient data references for deep reasoning, thus impairing its inference capability. On the other hand, when $k$=6, the outcome is essentially the same as when $k$=4. We think that once the number of prompting examples reaches a certain threshold, increasing $k$ further does not significantly stimulate deeper reasoning in the LLM, and also consumes more resources. Therefore, we choose $k$=4.
Table 7: Abalation study of the selection of $k$.
[…]
2 0.232 0.119 0.282 0.0009 0.835 0.797 0.0019 0.999
4 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
6 0.216 0.075 0.301 0.0004 0.968 0.939 0.0010 1.000
Effectiveness of the Multi‑Stage of CoT. To verify that a multi‑stage chain of thought is more effective than a single‑stage approach, we combined all sub‑problems from the multi‑stage process into a single question and performed a one‑stage refinement. The quantitative metrics, presented in Table. 8, show that the single‑stage strategy performs considerably worse than the multi‑stage variant. This disparity arises because layout generation follows a logical sequence: resolving sub‑problems step by step enhances layout rationality, whereas tackling multiple issues simultaneously is impractical and yields limited improvements.
Table 8: Abalation study of the effectiveness of multi-stage.
[…]
1 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
3 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
5 Conclusion
This paper focuses on unleashing the deep reasoning abilities of LLMs in layout generation tasks without requiring any training. To address the limitations of existing methods and enhance models’ performance in layout generation, we propose LayoutCoT, which comprises three key components: a Layout-aware RAG, a Coarse Layout Generator, and a Chain of Thought for Layout Refinement. Specifically, we select prompting examples through a carefully designed RAG approach to generate a coarse layout, then further decompose the layout task into multiple simpler subproblems. By integrating the RAG results, we refine the coarse layout. Our method activates the deep reasoning capabilities of LLMs in layout generation, achieving state-of-the-art results on five public datasets spanning three different types of layout tasks. These findings demonstrate that LayoutCoT is a highly versatile, data-efficient, and training-free approach that produces high-quality layouts while satisfying given constraints.