### v3

# LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation

Abstract
Conditional layout generation aims to automatically produce visually appealing and semantically coherent layouts under user-defined constraints. Generative models excel but require large task-specific datasets or costly fine-tuning, hampering scalability and real-world adoption. Training-free LLM methods, though cheaper, still suffer from limited reasoning depth and naive ranking strategies, leading to sub-optimal layouts. To this end, we introduce LayoutCoT, the first framework that explicitly unlocks the latent reasoning abilities of off-the-shelf LLMs for layout generation by uniting Layout-aware Retrieval-Augmented Generation (RAG) with Chain-of-Thought (CoT) reasoning. LayoutCoT first linearizes layouts into a LLM-friendly serialized form. Then it employs a layout-aware RAG module to retrieve relevant exemplars and then prompt the LLM to draft a coarse layout with the user’s constraints. Finally, we introduce a dedicated CoT module to iteratively refine this draft, reasoning step-by-step over spatial and semantic relations to produce layouts with markedly higher coherence and visual quality. On five public benchmarks covering three conditional layout generation tasks, LayoutCoT establishes new state-of-the-art results without fine-tuning. Remarkably, our CoT module enables a standard LLM to surpass specialized deep-reasoning models such as DeepSeek-R1, underscoring the promise of CoT-driven LLMs for scalable, training-free layout generation.
LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation
Hengyu Shithanks: Both authors contributed equally to this research. Names are ordered alphabetically., Junhao Su11footnotemark: 1, Tianyang Han, Junfeng Luo, Jialin Gaothanks: Corresponding author. Meituan {shihengyu02,sujunhao02,hantianyang,luojunfeng,gaojialin04}@meituan.com
Figure 1: Qualitative evaluation of the multi-stage LayoutCoT framework. Each successive stage systematically refines the spatial arrangement of layout elements, resulting in designs with enhanced coherence and rationality.
1 Introduction
Layout is a fundamental element in graphic design, determined by the deliberate and structured arrangement of visual components. Automated layout generation has recently emerged as an important research topic, aiming to significantly alleviate the workload of designers while addressing diverse user requirements Gupta et al. (2021); Jiang et al. (2022); Kikuchi et al. (2021); Kong et al. (2022); Lee et al. (2020); Li et al. (2020). Current automated layout generation tasks broadly fall into three categories: i) content-aware layouts, which integrate visual or textual content for contextually appropriate designs (e.g., aligning text boxes with salient image regions in posters); ii) constraint-explicit layouts, which strictly adhere to predefined constraints, commonly found in user interface (UI) design and document typesetting; iii) and text-to-layout tasks, which translate abstract textual instructions into spatially coherent layouts, crucial for cross-modal creative tools.
Early GAN and diffusion-based approaches Goodfellow et al. (2014); Ho et al. (2020) learned implicit design rules from extensive labeled datasets. Methods such as LayoutGAN Li et al. (2019) and LayoutDiffusion Zheng et al. (2023) have shown promising capabilities in generating plausible layouts, yet their reliance on substantial training data and domain-specific fine-tuning limits scalability and versatility. When faced with limited training data, purely generative methods Jiang et al. (2022); Hui et al. (2023); Inoue et al. (2023); Hsu et al. (2023); Lin et al. (2023b) often fail to effectively capture sparse data distributions, restricting generalization and necessitating separate models for distinct tasks. Furthermore, the demand for extensive labeled datasets makes training generative layout models costly and inefficient, highlighting the necessity for a more flexible and versatile layout generation paradigm.
With the advancement of large language models (LLMs) Brown et al. (2020); Imani et al. (2024); Achiam et al. (2023); Yang et al. (2022a, b), layout generation methods are shifting towards a training-free paradigm. Recent research has shown that LLMs inherently possess a certain understanding of layout principles, such as alignment and spatial coherence Imani et al. (2024); Chen et al. (2021). While leveraging this capability, approaches such as LayoutPrompter Lin et al. (2023b) utilize in-context learning combined with post-processing ranking mechanisms to guide LLMs in layout generation. However, this methodology demonstrates excessive focus on quantitative evaluation metrics, frequently resulting in structurally impractical bounding box configurations containing either disproportionately small elements or physically unfeasible arrangements. Moreover, when confronted with complex layout scenarios, the simplistic reasoning paradigms and elementary ranking systems inherent in such approaches tend to produce suboptimal compositions that violate both human visual perception principles and real-world application constraints.
To unleash the deep reasoning abilities of LLMs in layout generation, we must resolve two key issues: first, how to select suitable contexts as references for layout generation; and second, how to design a chain of thought that enables LLMs to engage in human-like deep reasoning during layout design. Therefore, we propose LayoutCoT, a novel approach that integrates a layout-aware RAG with Chain-of-Thought (CoT) in LLMs to address those above limitations. Specifically, LayoutCoT employs a layout-aware exemplar retrieval mechanism, dynamically selecting suitable reference contexts using layout similarity Otani et al. (2024), which better captures structural and semantic similarities compared to previous methods like LayoutPrompter Lin et al. (2023b). We further introduce a meticulously designed Chain-of-Thought (CoT) reasoning framework that refines the layout through three sequential stages, each addressing a distinct aspect of the design process. In the first stage, the LLM is prompted to evaluate each element from the specified categories and determine their initial positions, emphasizing logical structure and visual appeal without accounting for interactions among element sizes. In the second stage, the LLM is guided to optimize the sizes and positions of these elements, effectively resolving overlaps or potential crowding issues. Finally, in the third stage, fine-grained mathematical adjustments are applied to specific layout dimensions to avoid unreasonable arrangements—such as the occlusion of salient elements—and to ensure both aesthetic quality and practical viability.
We conducted comprehensive experiments across five public datasets (PKU Hsu et al. (2023), CGL Zhou et al. (2022a), RICO Liu et al. (2018), Publaynet Zhong et al. (2019), and WebUI Wu et al. (2023)), spanning three distinct layout generation tasks. Our results demonstrate that LayoutCoT consistently outperforms existing state-of-the-art methods without requiring any training or fine-tuning. Notably, our CoT module enables standard LLMs, such as GPT-4, to surpass the performance of specialized deep-reasoning models like DeepSeek-R1 Guo et al. (2025), underscoring the potential of our approach for practical layout generation tasks. The main contributions of our work are summarized as follows:
  • We introduce LayoutCoT, a method that effectively harnesses the deep reasoning capabilities of general-purpose LLMs through Chain-of-Thought prompting, significantly improving the performance and practicality of training-free layout generation.
  • LayoutCoT requires no task-specific training or fine-tuning, efficiently producing high-quality, visually coherent layouts by decomposing complex generation tasks into clear reasoning steps.
  • Extensive empirical validation demonstrates LayoutCoT’s superior performance and versatility across various tasks and datasets, highlighting its ability to elevate standard LLMs above specialized deep-reasoning counterparts.
Figure 2: Overview of the LayoutCoT framework. Our training-free approach initially employs a Layout-aware RAG to prompt the LLM for a coarse layout prediction, establishing a logical and visually coherent arrangement of elements. This is subsequently refined via a multi-stage Chain-of-Thought (CoT) module, which iteratively enhances the layout by resolving spatial conflicts and fine-tuning dimensions. The proposed framework is versatile and applicable to a wide array of layout generation tasks.
2 Related Work
Layout generation has emerged as an active area of research in recent years. It can be broadly divided into three categories: content-aware layout generation Hsu et al. (2023); Zheng et al. (2019), constraint-explicit layout generation Kikuchi et al. (2021); Kong et al. (2022); Lee et al. (2020), and text-to-layout generation Huang et al. (2021); Lin et al. (2023a). Content-aware layout generation aims to place layout elements on a given canvas in a visually pleasing manner while avoiding overlap with any background images on the canvas. Constraint-Explicit layout generation typically involves arranging predefined elements, or elements with specified dimensions, or even specified relationships among elements, on a blank interface in a reasonable configuration. Text-to-layout generation, meanwhile, focuses on creating layouts based on natural language descriptions, which is generally more challenging.
2.1 Generative layout generation
Early work in this field commonly utilized generative models for layout generation. Research based on Generative Adversarial Networks (GANs) Li et al. (2020); Hsu et al. (2023) and Variational Autoencoders (VAEs) Lee et al. (2020); Jyothi et al. (2019) pioneered the application of generative models to layout generation, achieving noticeable success in improving layout aesthetics through latent variable optimization. With the advent of the Transformer Vaswani et al. (2017) architecture, numerous studies Gupta et al. (2021); Kong et al. (2022); Huang et al. (2021); Lin et al. (2023a); Rahman et al. (2021); Hu et al. (2025); Han et al. (2025) emerged that adopted a sequence-to-sequence approach, further enhancing the quality of generated layouts. More recently, the proposal of diffusion Ho et al. (2020) models has led to additional methods Hui et al. (2023); Inoue et al. (2023) that bolster generative capabilities in scenarios with explicit layout constraints, offering further advancements in both quality and overall visual appeal. However, the existing generative layout methods are not capable of handling all types of layout generation tasks using a single model. Moreover, these approaches typically rely heavily on extensive datasets and require substantial training or fine-tuning, making practical applications challenging.
2.2 LLM-based Layout Generation
Large language models have demonstrated exceptional few-shot performance Brown et al. (2020); Achiam et al. (2023); Chowdhery et al. (2023); Touvron et al. (2023) across various natural language processing (NLP) tasks. Recent studies have shown that LLMs themselves can generate layouts in structured descriptive languages such as HTML Lin et al. (2023b); Tang et al. (2023). For instance, LayoutNUWA Tang et al. (2023) leverages Code Llama’s Roziere et al. (2023) code completion for layout generation, while PosterLlama Seol et al. (2024) fine-tunes models such as MiniGPT-4 Zhu et al. (2023) to further enhance code completion capabilities for layout generation tasks. However, these methods rely heavily on sophisticated fine-tuning strategies and are dependent on large amounts of data. In contrast, LayoutPrompter Lin et al. (2023b) and LayoutGPT Feng et al. (2023) utilize GPT-3 Brown et al. (2020) via in-context learning to achieve training-free layout generation, revealing that LLMs inherently possess the capacity for layout generation. Although these methods do not require extensive, painstaking training and can generalize across various tasks, the quality of their generated layouts remains a concern.
2.3 Retrieval-Augmented Generation and Chain of Thought
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a cutting-edge technique that combines retrieval and generation to enhance the model’s generative capabilities. By retrieving relevant examples or information Liu et al. (2024); Cheng et al. (2025), RAG introduces more comprehensive contextual data during the generation process, thereby improving the quality of the outputs. Content-aware graphic layout generation aims to automatically arrange visual elements based on given content, such as e-commerce product images. In this context, the Retrieval-Augmented Layout Transformer (RALF) Horita et al. (2024) leverages the retrieval of layout examples most similar to the input image to assist in layout design tasks. However, the RAG in RALF is constrained by its input paradigm, lacking plug-and-play characteristics and performing suboptimally across other tasks. Chain-of-Thought (CoT) Wei et al. (2022); Chen et al. (2025); Luo et al. (2025) is a method that enhances model reasoning capabilities through the simulation of human thought processes. CoT achieves this by progressively solving problems through phased reasoning, offering more interpretable and coherent solutions to complex tasks. Although CoT has demonstrated powerful reasoning advantages across numerous domains, it has not yet been applied to layout design tasks. We attempt to integrate RAG with CoT for layout generation tasks, utilizing RAG to provide rich contextual information, while employing CoT for phased reasoning and optimization. This synergistic approach aims to enhance the quality and coherence of the generated results, combining the strengths of both methodologies effectively.
Table 1: Dataset statistics. Training Set refers to the candidate pool for RAG example selection.
Dataset Tasks Training Set Test Set Element Types
PKU content-aware 9,974 905 3
CGL content-aware 38,510 1,647 5
RICO constraint-explicit 31,694 3,729 25
PubLayNet constraint-explicit 311,397 10,998 5
WebUI text-to-layout 3,835 487 10
3 Methods
Figure. 2 illustrates the overall structure of LayoutCoT, which consists of three main components: the Layout-aware RAG, the layout coarse generator, and the layout CoT module. Specifically, given the labels and bounding boxes (bboxes) of an input example layout, we use the Layout-aware RAG to select the top-k relevant layout examples as prompting examples. Next, the layout coarse generator takes these prompting examples and the given constraints to produce an initial coarse layout. Finally, both the prompting examples and this coarse layout are fed into the layout CoT module for fine-grained layout adjustments. Because the layout coarse generator lacks a detailed reasoning process, relying solely on the prompting examples and constraints may not fully satisfy the design requirements. Therefore, the layout CoT module breaks down the layout task into a multi-stage thought process; each stage refines the layout, ultimately yielding a more precise and aesthetically pleasing result.
Figure 3: Details of the CoT Module. We illustrate the overall conversational logic of the multi-stage CoT. Depending on the type of task, there are slight variations in the details to adapt to specific requirements. Some of LayoutCoT prompts can be found in the supplementary material.
3.1 Layout-aware RAG
We employ a Retrieval-Augmented Generation (RAG) mechanism to dynamically select relevant layout examples as prompts, thereby enhancing the flexibility and coherence of generation. Specifically, let $\mathcal{L}=\bigl\{(b_{i},c_{i})\bigr\}_{i=1}^{m}$ and $\hat{\mathcal{L}}=\bigl\{(\hat{b}_{j},\hat{c}_{j})\bigr\}_{j=1}^{n}$ denote two layouts, where each element consists of a bounding box $b_{i}$ (or $\hat{b}_{j}$) and an associated label $c_{i}$ (or $\hat{c}_{j}$). We use a dissimilarity function from LTSim Otani et al. (2024) D(,^)D($\mathcal{L}$$\hat{\mathcal{L}}$) based on a soft matching scheme:
$D(\mathcal{L},\hat{\mathcal{L}})=\min_{\Gamma\geq 0}\sum_{(i,j)\in\Omega}\Gamma_{i,j}\,\mu\bigl(b_{i},c_{i},\hat{b}_{j},\hat{c}_{j}\bigr)$ (1)
where $\Gamma_{i,j}$ represents a fractional assignment aligning element $(b_{i},c_{i})$ with $(\hat{b}_{j},\hat{c}_{j})$, and $\mu(\cdot)$ is a cost function that captures positional and label discrepancies. By solving this optimization, we obtain the minimal “transport cost” between the two layouts.
Next, we transform this cost-based measure into a layout-level similarity via an exponential mapping:
[…]
A lower cost D(,^)D($\mathcal{L}$$\hat{\mathcal{L}}$) implies higher layout similarity; we set the scale parameter to $1.0$ for simplicity Otani et al. (2024).
To retrieve the top-$K$ examples for a given layout $\mathcal{L}$, we consider a database $\mathbf{DB}=\{\hat{\mathcal{L}}_{1},\hat{\mathcal{L}}_{2},\dots,\hat{\mathcal{L}}_{N}\}$. We compute $\mathrm{LTSim}\bigl(\mathcal{L},\hat{\mathcal{L}}_{i}\bigr)$ for each candidate $\hat{\mathcal{L}}_{i}\in\mathbf{DB}$, and select the top-$K$ layouts that yield the highest scores:
[…]
We then collect these $\hat{\mathcal{L}}_{k}\in\mathcal{R}_{K}(\mathcal{L})$ as prompting examples in subsequent generation stages. By leveraging the soft matching mechanism in $\mathrm{LTSim}$, the Layout-aware RAG remains robust even when the input and candidate layouts differ in the number or types of elements.
Table 2: Quantitative comparison on Content-aware Layout Generation. We report results on both PKU (top) and CGL (bottom) datasets. LayoutCoT (Ours) consistently outperforms baselines across most metrics. Note that denotes variants using Gemini-2.5-pro, and denotes using DeepSeek-R1 without CoT.
Method Training-free Content Graphic
[…]
Dataset: PKU
CGL-GAN Zhou et al. (2022b) 0.219 0.175 0.226 0.0062 0.575 0.259 0.0605 0.707
DS-GAN Hsu et al. (2023) 0.209 0.217 0.229 0.0046 0.585 0.424 0.0260 0.903
RALF Horita et al. (2024) 0.171 0.150 0.246 0.0007 0.893 0.840 0.0180 0.998 0.981
LayoutPrompter Lin et al. (2023b) 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997 0.865
LayoutPrompter Lin et al. (2023b) 0.184 0.179 0.215 0.0025 0.799 0.751 0.0011 0.995 0.869
LayoutCoT 0.214 0.180 0.264 0.0016 0.925 0.849 0.0004 1.000 0.806
LayoutCoT (ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000 1.000
LayoutCoT (ours) 0.143 0.139 0.218 0.0011 0.821 0.770 0.0024 1.000 1.000
Dataset: CGL
CGL-GAN Zhou et al. (2022b) 0.489 0.268 0.147 0.0420 0.275 0.244 0.269 0.876
DS-GAN Hsu et al. (2023) 0.451 0.224 0.193 0.0485 0.370 0.301 0.075 0.893
RALF Horita et al. (2024) 0.336 0.197 0.247 0.0023 0.943 0.884 0.027 0.995 0.874
LayoutPrompter Lin et al. (2023b) 0.415 0.184 0.227 0.0091 0.338 0.317 0.0049 0.996 0.812
LayoutCoT 0.267 0.182 0.225 0.0018 0.515 0.501 0.0023 0.994 0.803
LayoutCoT (ours) 0.170 0.101 0.260 0.0005 0.958 0.958 0.0018 0.998 0.939
3.2 Coarse Layout Generation
After retrieving the top-$K$ examples via RAG (Section 3.1), we convert each layout $\hat{\mathcal{L}}_{k}\!\in\!\mathcal{R}_{K}(\mathcal{L})$ into a pre‑defined HTML snippet, the format definition of the HTML snippets is consistent with that of LayoutPrompter. Let the unified prompt set be
$\mathcal{S}\;=\;\bigl\{\mathrm{HTML}(\hat{\mathcal{L}}_{1}),\,\dots,\,\mathrm{HTML}(\hat{\mathcal{L}}_{K}),\;c\bigr\}$ (4)
where $c$ contains user‑defined constraints. Feeding $\mathcal{S}$ to the LLM yields $n$ candidate layouts in parallel:
$\bigl\{\mathcal{L}^{j}_{\mathrm{HTML}}\bigr\}_{j=1}^{n}=\text{LLM}(\mathcal{S})$ (5)
We score each candidate with a ranker $\tau(\cdot)$ Lin et al. (2023b) and select the highest‑scoring one as the coarse result:
$\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}\;=\;\arg\max_{\,1\leq j\leq n}\;\tau\bigl(\mathcal{L}^{j}_{\mathrm{HTML}}\bigr)$ (6)
This coarse layout preserves the user’s key structural constraints and serves as the starting point for refinement.
3.3 Chain of Thought for Layout Refinement
Given $\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}$ and the same constraint $c$, we apply a multi‑stage Chain‑of‑Thought strategy to iteratively improve the design. Let the task be decomposed into $i$ refinement sub‑problems $\{Q_{L}^{1},\dots,Q_{L}^{i}\}$. Denote $\mathcal{L}^{\mathrm{refine}_{0}}_{\mathrm{HTML}}=\mathcal{L}^{\mathrm{coarse}}_{\mathrm{HTML}}$. At stage $t$ ($1\!\leq\!t\!\leq\!3$) we query the LLM with the augmented prompt $\mathcal{S}\cup\{\mathcal{L}^{\mathrm{refine}_{t-1}}_{\mathrm{HTML}},\,Q_{L}^{t}\}$:
$\mathcal{L}^{\mathrm{refine}_{t}}_{\mathrm{HTML}}=\text{LLM}\Bigl(\mathcal{S}\cup\bigl\{\mathcal{L}^{\mathrm{refine}_{t-1}}_{\mathrm{HTML}},\,Q_{L}^{t}\bigr\}\Bigr),\,t=1,2,3$
The final refined layout $\mathcal{L}^{\mathrm{refine}_{3}}_{\mathrm{HTML}}$ is thus obtained after $3$ sequential reasoning steps, yielding a design that is both semantically coherent and visually pleasing.
4 Experiments
4.1 Experimental Setup
We conduct detailed experiments on three different types of tasks: content-aware, constraint-explicit, and text-to-layout, using five distinct datasets: PKU Hsu et al. (2023), CGL Zhou et al. (2022a), RICO Liu et al. (2018), Publaynet Zhong et al. (2019), and WebUI Wu et al. (2023). Since no official test set splits are provided for the PKU and CGL datasets, we used 9,974 labeled images in the PKU dataset as the training set and the database for the Layout-aware RAG module, with 905 unlabeled images serving as the test set. In the CGL dataset, we took 38,510 images as the training set and the database for the Layout-aware RAG module, and selected an additional 1,647 images (with no overlap with the training set) as the test set. For other datasets, we adopt the official data splits, as detailed in Table .1
4.2 Implement Details
In our experiments, we provide details on the choice of the $k$-value for the top-$k$ retrieval in the Layout-aware RAG. When generating the Coarse Layout, we follow the approach of LayoutPrompter Lin et al. (2023b) and set kk=10. For the CoT stage, we set kk=4. We also adopt LayoutPrompter’s setting of nn=10 for the number of generation times by LLMs when selecting the Coarse Layout. Since the ’Text-Davinci-003’ model used by LayoutPrompter is no longer available, we replaced the LLM with ’gpt-4’ in our comparative experiments. Additionally, on the PKU dataset, we conducted extra experiments using Gemini-2.5-pro, denoted as LayoutPrompter. In the LayoutCoT experiments, we used ’gpt-4’ as the LLM. To further validate the effectiveness of our CoT module, we provided LayoutCoT results for each experiment; LayoutCoT refers to the variant where the LLM is replaced by the reasoning model ’DeepSeek-R1’ and the CoT module is removed (only LayoutAware RAG and Coarse Layout Generation are used). For the PKU dataset, we also conducted experiments using Gemini-2.5-pro, denoted as LayoutCoT.
Figure 4: Qualitative Results for the Content-aware Layout Generation Task. LayoutPrompter and LayoutCoT tend to generate dense, small boxes in the upper-left corner, which is unreasonable. In contrast, LayoutCoT effectively corrects this error and achieves satisfactory results.
4.3 Evaluation Metrics
For the constraint-explicit layout generation and text-to-layout tasks, we utilize four standard metrics for evaluation: Alignment measures how well the elements within the layout align with each other. Overlap calculates the area of overlap between any two elements in the layout. Maximum IoU, mIoU assesses the maximum intersection over union between the generated layout and the actual layout, providing a measure of layout accuracy. Frechet Inception Distance (FID) quantifies the similarity between the distribution of generated layouts and the distribution of real layouts, using features extracted from a deep neural network trained on images. For the content-aware layout generation task, we reference the eight metrics used in Hsu et al. (2023) along with a newly designed metric $R_{e}$ that we introduce. The metric $R_{e}$ is used to assess the reasonableness of each element’s size within the layout. Let
$r_{i}=\frac{S^{i}}{S^{i}_{\text{train}}},\qquad d_{i}=\bigl|\ln r_{i}\bigr|,\qquad\tau=\ln(1.1)$
where $S^{i}$ is the predicted mean area of label $i$ and $S^{i}_{\text{train}}$ is its training-set reference. With a $\pm 10\%$ tolerance band $\tau$, we define
$\operatorname{score}_{i}=\exp\,\!\bigl(-\max(0,\,d_{i}-\tau)\bigr)$ (8)
Thus $\operatorname{score}_{i}=1$ when $0.9\leq r_{i}\leq 1.1$ and decays exponentially once the deviation exceeds $10\%$.
Aggregating over all $N$ labels yields
$R_{e}=\exp\Bigl(-\sqrt{1/N\sum\nolimits_{i=1}^{N}\bigl(\max(0,\,d_{i}-\tau)\bigr)^{2}}\Bigr)$ (9)
where $R_{e}\in(0,1]$; higher values indicate more reasonable element sizes.
4.4 Results on Content-Aware Layout Generation
The quantitative experimental results for both PKU and CGL datasets are summarized in Table 2.. As seen from the table, our method surpasses all the evaluation metrics used in content-aware layout generation tasks compared to both the training-based baseline methods and another training-free method, LayoutPrompter Lin et al. (2023b). Notably, our proposed CoT module enables "GPT-4" to even outperform the deep reasoning large language model "DeepSeek-R1" in this layout generation task. Moreover, our performance on the Overlap metric in the PKU dataset was suboptimal, prompting further visual analysis. As illustrated in Figure. 4, the very low scores on the Overlap metric for both LayoutPrompter and LayoutCoT result from numerous bad cases. In these cases, layout bounding boxes are extremely small and densely clustered in the upper-left corner, leading to low Overlap values but highly unaesthetic and impractical layout outcomes. Hence, we introduce the $R_{e}$ metric in Sec. 4.3 to measure the reasonableness of layout sizes, on which we have performed exceptionally well. Qualitative comparative cases are further provided in Figure. 4.
Table 3: Quantitative comparison with other methods on constraint-explicit layout generation tasks.
RICO PubLayNet
[…]
Gen-T BLT Kong et al. (2022) 0.216 25.63 0.150 0.983 0.140 38.68 0.036 0.196
LayoutFormer++ Jiang et al. (2022) 0.432 1.096 0.230 0.530 0.348 8.411 0.020 0.008
LayoutPrompter Lin et al. (2023b) 0.604 2.134 0.002 0.029 0.619 1.648 0.001 0.004
LayoutCoT 0.718 4.970 0.002 0.021 0.795 2.608 0.000 0.003
LayoutCoT (Ours) 0.719 2.189 0.001 0.022 0.772 1.227 0.000 0.002
Gen-TS BLT Kong et al. (2022) 0.604 0.951 0.181 0.660 0.428 7.914 0.021 0.419
LayoutFormer++ Jiang et al. (2022) 0.620 0.757 0.202 0.542 0.471 0.720 0.024 0.037
LayoutPrompter Lin et al. (2023b) 0.707 3.072 0.002 0.051 0.727 1.863 0.001 0.076
LayoutCoT 0.782 1.252 0.001 0.033 0.844 1.701 0.000 0.034
LayoutCoT (Ours) 0.795 3.506 0.000 0.031 0.847 1.430 0.000 0.004
Gen-R CLG-LO Kikuchi et al. (2021) 0.286 8.898 0.311 0.615 0.277 19.73 0.123 0.200
LayoutFormer++ Jiang et al. (2022) 0.424 5.972 0.332 0.537 0.353 4.954 0.025 0.076
LayoutPrompter Lin et al. (2023b) 0.583 2.249 0.001 0.044 0.627 2.224 0.001 0.011
LayoutCoT 0.724 2.399 0.001 0.030 0.779 1.853 0.000 0.004
LayoutCoT (Ours) 0.725 3.569 0.001 0.015 0.769 1.408 0.001 0.002
Comp. LayoutTrans Gupta et al. (2021) 0.363 6.679 0.194 0.478 0.077 14.76 0.019 0.001
LayoutFormer++ Jiang et al. (2022) 0.732 4.574 0.077 0.487 0.471 10.25 0.020 0.002
LayoutPrompter Lin et al. (2023b) 0.541 2.656 0.001 0.041 0.754 1.282 0.001 0.001
LayoutCoT 0.682 2.236 0.008 0.031 0.843 0.840 0.001 0.000
LayoutCoT (Ours) 0.716 2.112 0.001 0.012 0.770 0.531 0.000 0.001
Refine RUITE Rahman et al. (2021) 0.811 0.107 0.133 0.483 0.781 0.061 0.029 0.020
LayoutFormer++ Jiang et al. (2022) 0.816 0.032 0.123 0.489 0.785 0.086 0.024 0.006
LayoutPrompter Lin et al. (2023b) 0.874 0.225 0.001 0.110 0.756 0.143 0.003 0.001
LayoutCoT 0.869 0.876 0.001 0.071 0.801 0.265 0.001 0.000
LayoutCoT (Ours) 0.890 0.353 0.000 0.052 0.839 0.141 0.000 0.000
Figure 5: Qualitative Results for the Constraint-Explicit Layout Generation Task. The RICO dataset features a wide variety of label categories, making the completion task on RICO more challenging compared to other task types, we conduct visualizations based on the completion task of the RICO dataset. It is evident that LayoutCoT designs more rational layouts, with better element overlap and alignment compared to other methods.
4.5 Results on Constraint-Explicit Layout Generation
Table 3 presents the quantitative evaluation comparing LayoutCoT with several training-based baselines and the training-free method LayoutPrompter. From the metrics, it is evident that, even without any training, LayoutCoT achieves optimal performance on most indicators. The corresponding qualitative results in Figure 5 demonstrate that LayoutCoT exhibits outstanding layout control capabilities and generation quality. LayoutCoT not only effectively meets various input constraints–including element type, size, and relationship constraints–but also produces visually pleasing layouts by avoiding overlap between elements and enhancing overall alignment. These findings validate the effectiveness of LayoutCoT.
4.6 Results on Text-to-Layout
The quantitative metrics are presented in Table 4. Because text-to-layout is one of the most challenging tasks in layout generation, LayoutCoT falls short in the mIoU metric compared to the parse-then-place Lin et al. (2023a) method. However, it performs exceptionally well on the remaining three metrics. In many scenarios, LayoutCoT more effectively fulfills textual descriptions and excels in aesthetics, avoiding element overlap and maintaining element alignment.
Table 4: Quantitative comparison with baselines on text-to-layout. We compare our method with the baselines Mockup Huang et al. (2021), parse-then-place Lin et al. (2023a), and LayoutPrompter.
Methods mIoU $\uparrow$ FID $\downarrow$ Align. $\downarrow$ Overlap $\downarrow$
Mockup 0.193 37.012 0.0059 0.4348
parse-then-place 0.684 2.959 0.0008 0.1380
LayoutPrompter 0.174 4.773 0.0009 0.0107
LayoutCoT 0.164 1.018 0.0009 0.0068
LayoutCoT (Ours) 0.199 1.732 0.00006 0.0061
4.7 Ablation Studies
We validate the effectiveness of each component on the PKU dataset, as summarized in Table 5.
Impact of Modules. Table 5 (a) demonstrates that both the Layout-aware RAG and CoT modules individually improve performance over the baseline. Specifically, RAG enhances structural alignment (Align $0.0021\!\rightarrow\!0.0005$), while CoT effectively reduces occlusion. The full LayoutCoT framework achieves the best balance across all metrics, confirming the complementarity of retrieval context and reasoning capabilities. The slight rise in Overlap is a necessary trade-off to correct the baseline’s tendency to generate unrealistically small, crowded boxes.
Retrieval Strategy Analysis. Regarding the RAG module, Table 5 (b) shows that our LTSim-based retrieval significantly outperforms Random, mIoU, and DreamSim strategies by effectively capturing holistic layout similarity. Furthermore, we investigate the number of retrieved examples $k$ in Table 5 (c), finding that $k=4$ offers the optimal trade-off between providing sufficient context and minimizing noise.
CoT Design Validation. Finally, we analyze the reasoning mechanism. Table 5 (d) reveals that our multi-stage strategy markedly surpasses the single-stage approach, validating the necessity of decomposing complex layout tasks into logical steps. Moreover, Table 5 (e) confirms that our domain-specific manually designed CoT outperforms the generic Self-Refine mechanism in 7 out of 8 metrics, ensuring superior layout rationality.
Table 5: Comprehensive Ablation Studies on the PKU dataset. We systematically analyze the impact of (a) overall modules, (b) layout-aware retrieval strategies, (c) the number of retrieved examples ($k$), (d) reasoning stages, and (e) refinement methodologies. The best results are highlighted in bold.
[…]
(a) Overall Module Ablation
Baseline (No RAG/CoT) 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997
w/ RAG only 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
w/ CoT only 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
LayoutCoT (Full) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(b) Layout-aware RAG Strategy
No RAG 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
Random 0.299 0.184 0.241 0.0009 0.808 0.783 0.0023 0.997
mIoU 0.230 0.163 0.265 0.0003 0.951 0.894 0.0010 1.000
DreamSim 0.252 0.086 0.273 0.0004 0.972 0.932 0.0008 0.999
LTSim (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(c) Number of Retrieved Examples ($k$)
$k=2$ 0.232 0.119 0.282 0.0009 0.835 0.797 0.0019 0.999
$k=4$ (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
$k=6$ 0.216 0.075 0.301 0.0004 0.968 0.939 0.0010 1.000
(d) CoT Reasoning Stages
Single-Stage (1) 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
Multi-Stage (3) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(e) Refinement Method
Self-Refine 0.153 0.149 0.213 0.0012 0.868 0.806 0.0029 1.000
Manual CoT (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
5 Conclusion
This paper focuses on unleashing the deep reasoning abilities of LLMs in layout generation tasks without requiring any training. To address the limitations of existing methods and enhance models’ performance in layout generation, we propose LayoutCoT, which comprises three key components: a Layout-aware RAG, a Coarse Layout Generator, and a Chain of Thought for Layout Refinement. Specifically, we select prompting examples through a carefully designed RAG approach to generate a coarse layout, then further decompose the layout task into multiple simpler subproblems. By integrating the RAG results, we refine the coarse layout. Our method activates the deep reasoning capabilities of LLMs in layout generation, achieving state-of-the-art results on five public datasets spanning three different types of layout tasks. These findings demonstrate that LayoutCoT is a highly versatile, data-efficient, and training-free approach that produces high-quality layouts while satisfying given constraints.
6 Limitations
Our work focuses on unlocking the reasoning potential of LLMs, but this comes with trade-offs. The iterative nature of our CoT mechanism, while boosting performance, incurs higher computational overhead than non-autoregressive models. Additionally, since our method leverages in-context learning, its upper bound is influenced by the retrieval database; expanding the database with more diverse layouts would be necessary to handle extremely rare or complex constraints effectively.
Appendix A Prompt Examples of LayoutCoT
A.1 Content Aware Layout Generation
Stage 1: Initial Adjustment
System Prompt Role. You are an advanced layout design assistant, LayoutGPT. We will refine the layout in 3 separate steps. Each step is isolated, and you only have the references provided in the current context. Goal. Adjust element positions to achieve reasonable alignment without overlap. Procedure. 1. Adjust only text and logo elements (modify x/y position, size, and font-size). 2. Keep object and underlay locked (do not move or alter them). 3. Avoid overlap between text, logo, and object — even 1px overlap is disallowed. 4. Achieve reasonable alignment: for each logical group, logo should ideally appear above the corresponding text. 5. Return only updated HTML content with no extra explanation. Note. You may learn spatial relationships and layout logic from the reference layouts, but you must not directly copy element coordinates.
User Prompt Below are $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the initial layout needing refinement: {{CURRENT_HTML}}
Stage 2: Underlay Refinement
System Prompt Role. You are LayoutGPT. Continue refining the layout. Goal. Optimize the background elements based on the locked foreground. Procedure. 1. Keep object, text, and logo elements exactly as they were after Stage 1 (locked). 2. Adjust only underlay elements (background rectangles). 3. Avoid unintended overlap: underlay must not intersect with unrelated content. 4. Improve layout metrics: metrics_und_l: If underlay > target, maximize IoU with associated text/logo. metrics_und_s: If underlay $\leq$ target, ensure it fully contains the corresponding text/logo. 5. Return only updated HTML content.
User Prompt Below are the $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the layout after Stage 1: {{STAGE_1_HTML}} Please do a final check and minimal fixes as needed. Return only the final HTML.
Stage 3: Final Check
System Prompt Role. You are LayoutGPT. Perform the final check. Goal. Ensure zero unintended overlap and perfect alignment. Procedure. 1. Check for unintended overlap across all elements (text, logo, underlay, object). Ensure underlay only overlaps its own group. 2. Ensure alignment is preserved (e.g., logo above text). 3. Perform minimal fixes for small issues (e.g., slight 1-2px overlap, container issues). 4. Return only the final HTML result.
User Prompt Below are the $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the layout after Stage 2: {{STAGE_2_HTML}} Please do a final check and minimal fixes as needed. Return only the final HTML.
A.2 Constraint-Explicit Layout Generation (Gen-T)
We provide the prompt for the Gen-T task on the PubLayNet dataset as a representative example.
Stage 1: Main Element Placement
System Prompt Context. We have 5 element types: 1) text, 2) title, 3) list, 4) table, 5) figure. Goal. Initial placement of core text elements. Procedure. 1. Move/resize only title (2) and text (1). 2. Keep list (3), table (4), and figure (5) locked. 3. Minimize overlap between title & text and place them with good alignment. 4. Keep FID & maxIoU in mind. 5. Return only updated HTML.
User Prompt Here are $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Initial layout needing Stage 1 refinement: {{CURRENT_HTML}} Please adjust only title(2) & text(1) to reduce overlap and improve alignment.
Stage 2: Complex Element Placement
System Prompt Goal. Place complex elements while respecting Stage 1. Procedure. 1. Place/resize list (3), table (4), and figure (5). 2. Keep title (2) & text (1) from Stage 1 locked if possible. 3. Further reduce overlap among all 5 elements. 4. Refine alignment (e.g., align edges of tables with text). 5. Begin improving FID and maxIoU. 6. Return only updated HTML.
User Prompt $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Layout after Stage 1: {{STAGE_1_HTML}} Please refine list(3), table(4), figure(5) now. Return updated HTML.
Stage 3: Global Optimization
System Prompt Goal. Final thorough check for metrics optimization. Objectives. Alignment: Lower score implies better alignment. Overlap: Lower score implies minimal collisions. FID: Reduce unnatural layout styles. maxIoU: Match reference bounding arrangements. Procedure. 1. Enumerate each pair for overlap. 2. Check alignment lines or columns. 3. Evaluate if FID can be lowered by slight repositioning. 4. Shift/resize elements to improve IoU with references. 5. Provide brief chain-of-thought reasoning, then output final HTML.
User Prompt $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Layout after Stage 2: {{STAGE_2_HTML}} Perform your step-by-step reasoning to finalize alignment, overlap, FID, maxIoU. Return final HTML.
A.3 Text-to-Layout Generation
Stage 1: Parsing and Rough Layout
System Prompt Role. You are LayoutGPT. Goal. Parse text and establish initial structure. Procedure. 1. Parse text to identify required elements; add missing or remove extraneous ones. 2. Attempt strong overlap reduction from the start (moves or resizes). 3. Perform basic alignment ($\pm$10px). 4. Partially match reference bounding arrangement for mIoU if feasible. 5. Return updated HTML only.
User Prompt References (topk) total $LEN_{TOPK}$: {{REFERENCES_STR}} Current layout: {{CURRENT_HTML}} Text describing needed elements: {{TEXT_DESCRIPTION}} Please add/remove elements per text, reduce overlap, and do basic alignment. Return updated HTML.
Stage 2: Overlap Resolution
System Prompt Goal. Aggressive overlap reduction. Procedure. 1. Reduce Overlap to near zero (Top Priority). 2. Refine alignment lines ($\pm$5px). 3. Match bounding arrangement from references for higher mIoU. 4. Add or remove minor elements if beneficial. 5. Return updated HTML only.
User Prompt References: {{REFERENCES_STR}} Layout after Stage 1: {{STAGE_1_HTML}} Focus on Overlap $\to$ 0, align row/col $\pm$5px. Return updated HTML.
Stage 3: Semantic-Agnostic Refinement
System Prompt Goal. Final geometric polish. Procedure. 1. Temporarily ignore text meaning. Force Overlap = 0 among main elements. 2. Fix leftover collisions by moves/resizes. 3. Push bounding arrangement for better mIoU. 4. Fix FID style if it looks awkward. 5. Return updated HTML only.
User Prompt References: {{REFERENCES_STR}} Layout after Stage 2: {{STAGE_2_HTML}} Ignore text semantics, finalize Overlap=0, fix FID. Return updated HTML.